OpenStudy (anonymous):

Last one PLEAAASE HELP. I'm terrible in math Write the equation of the line in slope-intercept form with a slope of -5 and a y-intercept of -1.

jimthompson5910 (jim_thompson5910):

y = mx + b y = -5x + b ... plug in the given slope m = -5 y = -5x + (-1) ... plug in the given y-intercept b = -1 y = -5x - 1

jimthompson5910 (jim_thompson5910):

In general, y = mx + b describes any line with a slope of m and a y-intercept of b

OpenStudy (anonymous):

How do you get b?

OpenStudy (anonymous):

b is just the y-intercept

jimthompson5910 (jim_thompson5910):

b is the y-intercept

jimthompson5910 (jim_thompson5910):

in this case, the y-intercept is -1, so b = -1

OpenStudy (anonymous):

So x basically has no value?

jimthompson5910 (jim_thompson5910):

don't worry about x right now it takes on many values, but we don't know the value it takes on right now

OpenStudy (anonymous):

wouldn't the answer be y=-5x-1?

jimthompson5910 (jim_thompson5910):

that's what I wrote

OpenStudy (anonymous):

okay... didn't see that..lol

jimthompson5910 (jim_thompson5910):

lol that's fine

OpenStudy (anonymous):

Oh okay, I wish you wouldn't have done it for me. All you had to do was explain it.I kinda wanted to know how to do these lol But it's okay I am thankful for the help.

jimthompson5910 (jim_thompson5910):

oh sry about that

jimthompson5910 (jim_thompson5910):

ok here's another: what's the equation of the line with a slope of 1/2 and a y-intercept of 10

OpenStudy (anonymous):

y=mx+b y=1/2x+10 y=1/2 + 10 did I do it right?

jimthompson5910 (jim_thompson5910):

where did your x go?

jimthompson5910 (jim_thompson5910):

it should be \[\Large y = \frac{1}{2}x + 10\]

jimthompson5910 (jim_thompson5910):

assuming it's a typo, you got it right

OpenStudy (anonymous):

im sorry i meant to put it in there. but thank you! i actually understand something in math!

jimthompson5910 (jim_thompson5910):

yay that's great, and no worries happens to me all the time